OpenStudy (anonymous):

Integration of cotx(cosecx-1)e^x dx

OpenStudy (anonymous):

cotx(cosecx-1)e^x dx=\[\frac{ \cos x}{ \sin x }(\frac{ 1-\sin x }{ \sin x })e^x\]

OpenStudy (anonymous):

That is not the answer, am just putting the equality